# DreamRunner: Fine-Grained Compositional Story-to-Video Generation with Retrieval-Augmented Motion Adaptation

### v3

@@ Abstract
Storytelling video generation (SVG) aims to produce coherent and visually rich multi-scene videos that follow a structured narrative. Existing methods primarily employ LLM for high-level planning to decompose a story into scene-level descriptions, which are then independently generated and stitched together. However, these approaches struggle with generating high-quality videos aligned with the complex single-scene description, as visualizing such complex description involves coherent composition of multiple characters and events, complex motion synthesis and muti-character customization. To address these challenges, we propose DreamRunner, a novel story-to-video generation method: First, we structure the input script using a large language model (LLM) to facilitate both coarse-grained scene planning as well as fine-grained object-level layout planning. Next, DreamRunner presents retrieval-augmented test-time adaptation to capture target motion priors for objects in each scene, supporting diverse motion customization based on retrieved videos, thus facilitating the generation of new videos with complex, scripted motions. Lastly, we propose a novel spatial-temporal region-based 3D attention and prior injection module SR3AI for fine-grained object-motion binding and frame-by-frame spatial-temporal semantic control. We compare DreamRunner with various SVG baselines, demonstrating state-of-the-art performance in character consistency, text alignment, and smooth transitions. Additionally, DreamRunner exhibits strong fine-grained condition-following ability in compositional text-to-video generation, significantly outperforming baselines on T2V-ComBench. Finally, we validate DreamRunner’s robust ability to generate multi-object interactions with qualitative examples.
Figure 1: Overall pipeline for DreamRunner. (1) plan generation stage: we employ an LLM to craft a hierarchical video plan (i.e., “High-Level Plan” and “Fine-Grained Plan”) from a user-provided generic story narration. (2.1) motion retrieval and prior learning stage: we retrieve videos relevant to the desired motions from a video database for learning the motion prior through test-time fine-tuning. (2.2) subject prior learning stage: we use reference images for learning the subject prior through test-time fine-tuning. (3) video generation with region-based diffusion stage: we equipt diffusion model with a novel spatial-temporal region-based 3D attention and prior injection module (i.e., SR3AI) for video generation with fine-grained control.
@@ 1 Introduction
Advancing storytelling video generation (SVG) is crucial for real-world video generation applications, enabling the creation of rich, immersive narratives with multiple realistic scenes, characters, and interactive events. Unlike existing short-form video generation approaches [52, 20, 73, 11, 56, 15, 28, 62, 45, 71, 13, 8], these models allow characters and objects to evolve across scenes, enhancing the coherence of generated content to align more closely with human storytelling. Such capabilities hold vast potential in media, gaming, and interactive storytelling.
@@
Existing SVG methods [18, 81, 40, 78, 17, 74] primarily employ high-level planning with a large language model (LLM), breaking down a story into multiple key scene descriptions. Each scene is then generated independently as a separate video and later stitched together to form a complete long-form storytelling video. Generating high-quality single-scene video exhibits three key challenges: 1) Coherent composition: As a highly complex textual form, a story, even at the single-scene level (e.g. “a woman in red dress on the left and a man in a tailcoat on the right is walking towards each other, they meet in the middle and start ballroom dancing in the room”), typically involves multiple objects/characters with distinct motion trajectories, attributes, and sequentially occurring events, all of which must be coherently composed in the generated video. 2) Complex motion synthesis: The complex scene descriptions often feature intricate character motions (e.g. “ballroom dancing”) that are difficult to generate from the base text-to-video (T2V) models. 3) Multi-character customization: These descriptions usually involve multiple characters with pre-defined reference images (e.g. for both the woman and the man), requiring the T2V model to maintain their visual appearance consistent with the references. However, recent SVG methods often feed single-scene descriptions directly as textual conditions to the T2V model with limited constraints, resulting in suboptimal fidelity, missed events/objects, unclear motions, and support for only single-character customization.
@@
To address the aforementioned challenges, we propose DreamRunner, a novel SVG framework for enhancing the fine-grained alignment between scene descriptions and generated videos. The core idea is that, beyond high-level planning, we also leverage the LLM for compositional reasoning—decomposing complex scene descriptions into a detailed, frame-by-frame layout plan of multiple entities with sequential motions/events—followed by region-based attention to ensure plan adherence for coherent composition, Additionally, we employ a retrieval-augmented prior learning approach for complex motion synthesis while injecting learned priors only to its corresponding regions for multi-character customization. Specifically, DreamRunner presents three essential processes in the framework: (1) Dual-Level Video Plan Generation, (2) Motion Retrieval and Subject/Motion Prior Learning, and (3) Spatial-Temporal Region-Based 3D Attention and Prior Injection (SR3AI). In (1) plan generation stage, given a user-provided story narration (e.g. “write a story of the witch and her cat’s one day”), we employ an LLM for hierarchical planning: first generate a high-level plan with character-driven, motion-rich event descriptions across scenes, then decompose the scene descriptions into detailed, entity-specific frame-level layout plans within each scene The generated frame-level plan serves as the fine-grained guidance for T2V. In (2) prior learning stage, we learn both subject and motion priors to improve character consistency and motion fidelity. We learn customized subject priors [50] on character reference images to adapt the model to specific appearances. Then we frame complex motion synthesis as a customization task akin to character customization to learn motion priors, aiming to capture specific visual patterns of relevant motions. We propose a novel automatic retrieval pipeline that selects motion-relevant videos from a large-scale video database [58] as references. We then apply test-time fine-tuning [75] to learn customized motion priors. Note that unlike most motion customization methods that use a single prompt as textual condition (e.g.“A person is dancing”) for all reference videos, we propose using per-video prompts to improve motion specificity and customization performance. Both motion and character priors are learned by updating only parameter-efficient modules [22] within specific layers of DiT [44]. In (3) video generation, we introduce SR3AI, a novel spatial-temporal region-based 3D attention and prior injection module that enables fine-grained control without additional training. Unlike previous work only supports spatial control [37, 36, 66, 24] or temporal control [7], SR3AI utilizes the frame-level layout plan as guidance, allows spatial-temporal controls to generate temporally sequential events/motions, multiple objects with specific attributes, movement trajectories, spatial relationships, etc. To achieve this, we first encode multiple conditions from the fine-grained plan. Then, within SR3AI, we compute the corresponding visual latents in latent space for each condition based on the spatial-temporal layouts, ensuring each condition attends only to its designated spatial-temporal latent region via masking attention to unrelated areas. This spatial-temporal region-based attention provides precise control and coherent composition over multiple objects and motions/events. Additionally, we extend this region-based design to inject learned character and motion priors directly into their corresponding regions within the diffusion model, seamlessly integrating multiple priors while avoiding conflicts and achieving multi-character multi-motion customization.
@@
We validate the effectiveness of DreamRunner on two tasks: story-to-video generation and compositional text-to-video generation. For story-to-video generation, we collect a story dataset, DreamStorySet, and compare DreamRunner with prior SoTA methods (i.e., VideoDirectGPT [37] and VLogger [81]). DreamRunner achieves a 13.1% relative improvement in character consistency (CLIP + DINO score) and an 8.56% relative gain in text-following capability (CLIP + ViCLIP score) over these approaches. Additionally, it enhances sequential event generation within a single scene, achieving a 27.2% relative improvement in DINO score for smooth multi-event transitions, underscoring its effectiveness. In compositional text-to-video generation, DreamRunner outperforms baseline methods on T2V-CompBench [54] across all dimensions, demonstrating its strength in compositional generation. Notably, despite being based on open-source models [68], DreamRunner achieves the highest scores in dynamic attribute binding and object interaction, along with comparable results in spatial relationships and motion binding to closed-source models. This highlights the potential of our approach to bridge the performance gap between open- and closed-source models. Lastly, we provide qualitative examples to showcase the effectiveness of DreamRunner in coherent composition and multi-character customization, etc. In summary, our main contributions include:
@@
  • A spatiotemporal region-based attention module for coherent composition of multiple objects and sequential events, along with a region-based LoRA injection design for multi-character, multi-motion customization.
@@ 2 Related Work
Storytelling Video Generation. Storytelling video generation [18, 81, 40, 78, 17, 74, 39] aims to create multi-scene videos based on given scripts. VideoDirectorGPT [37] and Vlogger [81] use LLMs for high-level planning, decomposing scripts into multi-scene conditions, and generating videos scene by scene. Animate-A-Story [18] further improves motion control by using retrieval-augmented videos as depth conditions. Recently, DreamStory [17] and MovieDreamer [74] generate keyframes using text-to-image models, which are then animated with image-to-video models [9] to create a coherent flow. Customization methods [50, 14, 29, 53, 43, 77, 69, 60, 32] are also involved to keep the consistency of reference images. Unlike these approaches, our work is designed to address the video-centric challenge of generating multi-characters, motion-rich videos with smooth, natural transitions between different actions using dual-level LLM planning.
Compositional Generation. Recent advances in diffusion models have enhanced compositional text-to-video generation by improving coherence, semantic alignment, and user control. Several methods leverage LLMs for scene planning [36, 37, 13, 76, 46], while others employ regional masks for multi-object control [24, 55, 66, 70, 61] or frame-level semantic conditioning [7, 65]. Additionally, LoRA-based compositional techniques integrate diverse concepts within the generation process [67, 33, 16, 79]. However, these approaches do not explicitly bind objects to their corresponding actions/events spatial-temporally. Our method ensures fine-grained control over both objects and motions, maintaining a cohesive object-action link throughout the video.
Motion Customization. Motion customization is a key challenge in video generation. One approach focuses on pixel-level motion learning for video editing, aiming to replicate low-level motions from reference images while maintaining frame consistency [63, 72, 48, 25, 38]. Another emphasizes learning generic motion priors, such as human or camera movements, from curated video collections [75, 64, 80, 59], capturing high-level semantic actions for realistic motion synthesis. Most methods use test-time fine-tuning with motion LoRAs or adapters for adaptation. Our approach follows this pipeline but improves it by learning motion priors from retrieved videos in large-scale databases, enabling more diverse and contextually relevant motion customization. Additionally, we replace the single-prompt condition with per-video detailed prompts, enhancing motion specificity and generation quality.
@@ 3 Methodology
Task Setup. Storytelling video generation focuses on creating multi-scene, character-driven videos based on a given topic. The characters are defined by reference images (e.g., images of a witch), and the topic is presented as an instructional prompt (e.g., "witch’s one day"). The generated videos should align with the given topic and accurately reflect the characteristics and behavior of the characters.
Method Overview. Our approach employs a hierarchical system where an LLM generates event-based scripts across multiple scenes, followed by detailed plans specifying the layout and motion transitions of key objects per scene (Sec. 3.1). A video diffusion model then synthesizes each scene step by step. We train motion priors from retrieval videos aligned with the LLM-generated plans, sourced from a large-scale video-language database, and character priors using the reference images (Sec. 3.2). Finally, we inject these priors and detailed plans into the video generation process in a zero-shot manner using our spatial-temporal regional diffusion module SR3AI (Sec. 3.3).
Base Generation Model. We leverages CogVideoX [68] as the foundational text-to-video backbone. CogVideoX employs a DiT-based architecture that integrates full 3D attention, and generates 6-second videos at 8 fps conditioned on input text. In our method, we extend CogVideoX by training character and motion priors in distinct layers (see Sec. 3.2) and by modifying its 3D attention (see Sec. 3.3) for better motion and character binding.
@@ 3.1 Generating Dual-Level Plans with LLMs
Story-Level Croase-Grained Planning. We use an LLM to generate a sequence of scene descriptions. Specifically, given the task requirements, a single in-context example, and the specified story topic, we prompt GPT-4o [41] to generate 6$\sim$8 descriptions that focus on character-driven, motion-rich events. Each description is in the format of scene, motions, narrations, where we first define the related motions then generate corresponding event narration. This description list serves as a high-level catalog that directs the story’s progression across scenes, ensuring that each event is consistent with the narrative flow of the video.
Scene-Level Fine-Grained Planning. After generating a list of single-scene descriptions with narrations, we create detailed, entity-level plans for each latent frame. Each plan consists of an overall background description followed by entity-specific details for each latent frame. As shown in the yellow Frame-Level Plan box at the top of Fig. 2, the background provides a global scene description (e.g., “A large garden"), formatted as Background: background description. Entity-level details specify each entity’s description, motion (e.g., "A [v1] witch is walking among the moonlit flowers in her garden"), and bounding box layout, formatted as: Frame: [entity name, entity motion, entity description], [x0,y0,x1,y1]. Here, [x0,y0,x1,y1] denotes the top-left and bottom-right corners of the bounding box, with coordinates normalized to $[0,1]$. Entities without motion are labeled "none". Each scene includes plans for six key frames, with each frame guiding one second of video generation (we interpolate key frames to match the #frames of visual latents), resulting in a six-second output using CogVideoX. Detailed prompt templates for both levels’ planning are in Appendix H.
@@ 3.2 Motion Retrieval and Prior Learning
Retrieving Motion-Related Videos from Database. We employ a retrieval-augmented approach to fine-tune motion priors at test time, enhancing the model’s capacity to generate complex and diverse motions. Based on motion descriptions generated from the LLM planning, we retrieve relevant videos from a large-scale video database [58]. Our retrieval pipeline first uses BM25 [49] for initial text-based retrieval, followed by attribute-based filtering and clip segmentation via object tracking [26]. We then compute semantic similarity scores using CLIP [47] and ViCLIP [58] to refine the selection, ensuring high-quality motion-aligned videos (see Appendix A.1 for details). By following this process, we retrieve $4\sim 20$ video clips per motion, which are then used as reference videos for learning motion priors.
Motion Prior Training. We follow recent motion customization methods [75] to apply test-time fine-tuning for learning specific motion priors. Typically, reference videos are used to fine-tune a video diffusion model [56, 1] with LoRA, injecting temporal LoRAs into temporal attention layers to capture motion patterns (e.g., jumping) and per-video-specific LoRAs into spatial layers to capture unique video characteristics. Only temporal LoRAs are used during inference. Since our approach is based on CogVideoX [68], which employs 3D full attention instead of separate spatial and temporal attention, we manually designate even layers as “spatial” and odd layers as “temporal” to separate the learning of spatial and temporal LoRAs. We train LoRAs on top-ranked retrieved videos with all backbone parameters frozen, optimizing both a standard diffusion loss for frame reconstruction and an appearance-debiased temporal loss to focus on motion-specific learning. Loss details are in Appendix A.2. Note that unlike previous methods using a single-prompt condition for all the retrieved videos, we utilize video caption provided from the video database as per-video prompt. This helps the model implicitly separate motion-unrelated backgrounds, appearances, etc, allowing it to focus on motion-specific patterns. Note that when injecting spatial/temporal LoRAs to target transformer layers, we only inject them to the linear q, k, v projection in each block, and the final feedforward layers.
@@
VideoDirectorGPT [37] 54.3 9.5 23.7 21.7 22.4 22.5 63.5 42.3 60.3 94.3
VLogger [81] 62.5 41.3 23.5 23.1 22.5 22.2 73.6 43.4 61.2 96.2
DreamRunner (Ours) 70.7 (+13.1%) 55.1 (+33.4%) 24.7 (+5.11%) 23.7 (+2.60%) 24.2 (+7.56%) 24.1 (+8.56%) 93.6 (+27.2%) 55.4 (+27.6%) 62.1 (+1.47%) 98.1 (+1.98%)
Table 1: Evaluation of story-to-video generation on DreamStorySet. We compare our model with VideoDirectorGPT [37] and VLogger [81] on character consistency (via CLIP and DINO scores), text instructions following and full prompt adherence (via CLIP and ViCLIP scores), and event transitions smoothness (via DINO score). Our relative improvement over VLogger is highlighted in blue.

### v4

@@ Abstract
Storytelling video generation (SVG) aims to produce coherent and visually rich multi-scene videos that follow a structured narrative. Existing methods primarily employ LLM for high-level planning to decompose a story into scene-level descriptions, which are then independently generated and stitched together. However, these approaches struggle with generating high-quality videos aligned with the complex single-scene description, as visualizing such complex description involves coherent composition of multiple objects/events, complex motion synthesis and character customization with sequential motions. To address these challenges, we propose DreamRunner, a novel story-to-video generation method: First, we structure the input script using a large language model (LLM) to facilitate both coarse-grained scene planning as well as fine-grained object-level layout planning. Next, DreamRunner presents retrieval-augmented test-time adaptation to capture target motion priors for objects in each scene, supporting diverse motion customization based on retrieved videos, thus facilitating the generation of new videos with complex, scripted motions. Lastly, we propose a novel spatial-temporal region-based 3D attention and prior injection module SR3AI for fine-grained object-motion binding and frame-by-frame spatial-temporal semantic control. We compare DreamRunner with various SVG baselines, demonstrating state-of-the-art performance in character consistency, text alignment, and smooth transitions. Additionally, DreamRunner exhibits strong fine-grained condition-following ability in compositional text-to-video generation, significantly outperforming baselines on T2V-ComBench. Finally, we domonstrate DreamRunner’s ability to generate multi-character interactions with qualitative examples.
Figure 1: Overall pipeline for DreamRunner. (1) plan generation stage: we employ an LLM to craft a hierarchical video plan (i.elet@tokeneonedot, “High-Level Plan” and “Fine-Grained Plan”) from a user-provided generic story narration. (2.1) motion retrieval and prior learning stage: we retrieve videos relevant to the desired motions from a video database for learning the motion prior through test-time fine-tuning. (2.2) subject prior learning stage: we use reference images for learning the subject prior through test-time fine-tuning. (3) video generation with region-based diffusion stage: we equip diffusion model with a novel spatial-temporal region-based 3D attention and prior injection module (i.elet@tokeneonedot, SR3AI) for video generation with fine-grained control.
@@ 1 Introduction
Advancing storytelling video generation (SVG) is crucial for real-world video generation applications, enabling the creation of rich, immersive narratives with multiple realistic scenes, characters, and interactive events. Unlike existing short-form video generation approaches (singer2022make; hong2022cogvideo; zhang2023controlvideo; chen2023control; wang2023modelscope; girdhar2023emu; khachatryan2023text2video; weng2024art; qing2024hierarchical; zhang2024show; fei2024dysen; bar2024lumiere), these models allow characters and objects to evolve across scenes, enhancing the coherence of generated content to align more closely with human storytelling. Such capabilities hold vast potential in media, gaming, and interactive storytelling.
@@
Existing SVG methods (he2023animate; zhuang2024vlogger; oh2025mevg; zheng2024temporalstory; he2024dreamstory; zhao2024moviedreamer) primarily employ high-level planning with a large language model (LLM), breaking down a story into multiple key scene descriptions. Each scene is then generated independently as a separate video and later stitched together to form a complete long-form storytelling video. Generating high-quality single-scene video exhibits three key challenges: 1) Coherent composition: As a highly complex textual form, a story, even at the single-scene level (e.g. “Lucy on the left and a man on the right is walking towards each other, they meet in the middle and start ballroom dancing”), typically involves multiple objects/characters with distinct motion trajectories, attributes, and sequentially occurring events, all of which must be coherently composed in the generated video. 2) Complex motion synthesis: The complex scene descriptions often feature intricate character motions (e.g. “ballroom dancing”) that are difficult to generate from the base text-to-video (T2V) models. 3) Character customization with sequential events: These descriptions usually involve characters with pre-defined reference images (e.g. Lucy), with sequential motions (e.g. walking to ballroom dancing), making it challenging to maintain both temporal coherence and visual consistency with the character. However, recent SVG methods often feed single-scene descriptions directly as textual conditions to the T2V model with limited constraints, resulting in suboptimal fidelity, missed events/objects, unclear motions, etc.
@@
To address the above challenges, we propose DreamRunner, a novel SVG framework that enhances fine-grained alignment between scene descriptions and generated videos. Beyond high-level planning, DreamRunner uses LLM-based compositional reasoning to decompose complex scenes into frame-by-frame layout plans of multiple entities with sequential motions/events, followed by region-based attention for coherent composition. For complex motion synthesis, we adopt retrieval-augmented prior learning, injecting priors only into relevant regions to support character customization with sequential motions. Specifically, DreamRunner presents three essential processes in the framework: (1) Dual-Level Video Plan Generation, (2) Motion Retrieval and Subject/Motion Prior Learning, and (3) Spatial-Temporal Region-Based 3D Attention and Prior Injection (SR3AI). In (1) plan generation stage, given a user-provided story narration (e.g. “write a story of the witch and her cat’s one day”), we employ an LLM for hierarchical planning: first generate a high-level plan with character-driven, motion-rich event descriptions across scenes, then decompose the scene descriptions into detailed, entity-specific frame-level layout plans within each scene. The generated frame-level plan serves as the fine-grained guidance for T2V. In (2) prior learning stage, we learn both subject and motion priors to enhance character consistency and motion fidelity. Subject priors are learned from character reference images using customization techniques (ruiz2023dreambooth) to adapt the model to specific appearances. Then we treat complex motion synthesis as a customization problem and learn motion priors to capture the visual patterns of target motions. To this end, we introduce an automatic retrieval pipeline that selects motion-relevant videos from a large-scale dataset (wang2023internvid) as references. We then apply test-time fine-tuning (zhao2023motiondirector) to learn customized motion priors. We use per-video prompts—rather than a shared one as in prior methods—to improve motion specificity, and learn both priors via LoRA-based tuning (hu2021lora) on specific layers of DiT (peebles2023scalable). In (3) video generation stage, we introduce SR3AI, a novel spatial-temporal region-based 3D attention and prior injection module that enables fine-grained control without additional training. Unlike prior methods that support only spatial (lin2023videodirectorgpt; lianllm; dav24; jain2024peekaboo) or temporal (bansal2024talc) control, SR3AI leverages frame-level layout plans to enable spatial-temporal control over sequential events, object attributes, trajectories, and spatial relationships. We first encode multiple conditions from the fine-grained plan. SR3AI then computes visual latents for each condition based on its spatial-temporal layout and enforces attention masking so that each condition attends only to its designated region. This ensures precise control and coherent composition of multiple objects and motions. Moreover, we extend this region-based design to inject learned character and motion priors into their corresponding regions in the diffusion model, enabling coherent character and motion customization.
@@
We validate the effectiveness of DreamRunner on two tasks: story-to-video generation and compositional text-to-video generation. For SVG, we collect a story dataset, DreamStorySet, and compare DreamRunner with SoTA methods (VideoDirectGPT (lin2023videodirectorgpt) and VLogger (zhuang2024vlogger)). DreamRunner achieves a 13.1% relative improvement in character consistency score and an 8.56% gain in text alignment score. It also improves sequential event generation within a single scene, with a 27.2% boost for smoother multi-event transitions. Qualitative results further show strong generalization to multi-character settings. In compositional T2V generation, DreamRunner outperforms baseline methods on T2V-CompBench (sun2024t2v) across all dimensions, demonstrating its strength in compositional generation. Notably, despite being based on open-source models (yang2024cogvideox), DreamRunner achieves the highest scores in dynamic attribute binding and object interaction, along with comparable results in spatial relationships and motion binding to closed-source models, showing our method’s potential to bridge the performance gap between open- and closed-source models. In summary, our main contributions include:
@@
  • A spatiotemporal region-based attention module for coherent composition of multiple objects and sequential events, along with a region-based LoRA injection design for character and sequential motion customization.
@@ 2 Related Work
Storytelling Video Generation. Storytelling video generation (oh2025mevg; zheng2024temporalstory; long2024videostudio) aims to produce multi-scene videos from input scripts. Existing approaches use either high-level LLM planning for step-by-step decomposition and generation (zhuang2024vlogger; he2023animate; lin2023videodirectorgpt)or keyframe generation with text-to-image models followed by video animation (he2024dreamstory; zhao2024moviedreamer; chai2023stablevideo). Reference-based customization methods (ruiz2023dreambooth; TI23; custom_diffusion; StyleDrop; CAT24; StyleBreeder; ipadapter; ELITE; BLIPDiffusion23; chen2025multi; huang2025conceptmaster; liu2025phantom) to preserve character identity across scenes are also adopted. Our work targets the video-centric challenge of generating multi-character, motion-rich videos with smooth, natural transitions.
Compositional Generation. Recent advances in diffusion models have enhanced compositional T2V generation by improving coherence, semantic alignment, and user control. Several methods leverage LLMs for scene planning (lianllm; lin2023videodirectorgpt; fei2024dysen; zheng2023layoutdiffusion; qu2023layoutllm), while others employ regional masks for multi-object control (jain2024peekaboo; videoteris; dav24; yu2024zero; wei2024dreamvideo) or frame-level semantic conditioning (bansal2024talc; xing2024aid). Additionally, LoRA-based compositional techniques integrate diverse concepts within the generation process (yang2024loracomposer; li2024selma; gu2024mix; zhong2024multi). However, these approaches do not explicitly bind objects to their corresponding actions/events spatial-temporally. Our method ensures fine-grained control over both objects and motions, maintaining a cohesive object-action link throughout the video.
Motion Customization. Motion customization remains a key challenge in video generation. One line of work focuses on pixel-level motion learning for video editing, aiming to replicate fine-grained motions from reference frames while preserving temporal consistency (wu2023tune; zhang2023motioncrafter; ren2024customize; jeong2023vmc; lin2024ctrl). Another learns generic motion priors—e.g., human or camera movements—from curated datasets (zhao2023motiondirector; wu2023lamp; zhu2024echoreel; wei2023dreamvideo), capturing high-level semantics for realistic motion synthesis. Most approaches rely on test-time fine-tuning with motion LoRAs or adapters. Our method builds on this pipeline by retrieving motion-relevant videos from large-scale databases for more diverse and context-aware motion priors. We also replace the standard single-prompt conditioning with per-video detailed prompts to improve motion specificity and generation quality.
@@ 3 Methodology
Task Setup. Storytelling Video Generation focuses on creating multi-scene, character-driven videos based on a given topic. The characters are defined by reference images (e.g. images of a witch), and the topic is presented as an instructional prompt (e.g. "witch’s one day"). The generated videos should align with the given topic and accurately reflect the characteristics and behavior of the characters.
Method Overview. Our approach employs a hierarchical system where an LLM generates event-based scripts across multiple scenes, followed by detailed plans specifying the layout and motion transitions of key objects per scene (Section 3.1). A video diffusion model then synthesizes each scene step by step. We train motion priors from retrieval videos aligned with the LLM-generated plans, sourced from a large-scale video-language database, and character priors using the reference images (Section 3.2). Finally, we inject these priors and detailed plans into the video generation process in a zero-shot manner using our spatial-temporal regional diffusion module SR3AI (Section 3.3).
Base Generation Model. We leverages CogVideoX-2B (yang2024cogvideox) as the base text-to-video model. CogVideoX-2B employs a DiT-based architecture that integrates full 3D attention, and generates 6-second videos at 8 fps conditioned on input text. In our method, we extend CogVideoX-2B by training character and motion priors in distinct layers (see Sec. 3.2) and by modifying its 3D attention (see Sec. 3.3) for better motion and character binding.
@@ 3.1 Generating Dual-Level Plans with LLMs
[…]
Scene-Level Fine-Grained Planning. After generating a list of single-scene descriptions with narrations, we create detailed, entity-level plans for each latent frame. Each plan consists of an overall background description followed by entity-specific details for each latent frame. As shown in the yellow Frame-Level Plan box at the top of Figure 2, the background provides a global scene description (e.g., “A large garden"), formatted as Background: background description. Entity-level details specify each entity’s description, motion (e.g., "A [v1] witch is walking among the moonlit flowers in her garden"), and bounding box layout, formatted as: Frame: [entity name, entity motion, entity description], [x0,y0,x1,y1]. Here, [x0,y0,x1,y1] denotes the top-left and bottom-right corners of the bounding box, with coordinates normalized to $[0,1]$. Entities without motion are labeled "none". When bounding boxes overlap, we prompt the LLM to generate a unified caption that integrates the descriptions of all entities within the overlapping region. Each scene includes plans for six key frames, with each frame guiding one second of video generation (we interpolate key frames to match the #frames of visual latents), resulting in a six-second output using CogVideoX. Detailed prompt templates for both levels’ planning are in Appendix I.
@@ 3.2 Motion Retrieval and Prior Learning
Retrieving Motion-Related Videos from Database. We employ a retrieval-augmented approach to fine-tune motion priors at test time for complex motion synthesis. Based on motion descriptions generated from the LLM planning, we retrieve relevant videos from a large-scale video database (wang2023internvid). Our retrieval pipeline first uses BM25 (robertson2009probabilistic) for initial text-based retrieval, followed by attribute-based filtering and clip segmentation via object tracking (Jocher_YOLOv5_by_Ultralytics_2020). We then compute semantic similarity scores using CLIP (radford2021learning) and ViCLIP (wang2023internvid) to refine the selection, ensuring high-quality motion-aligned videos (see Appendix A.1 for details). By following this process, we retrieve $4\sim 20$ video clips per motion, which are then used as reference videos for learning motion priors.
Motion Prior Training. We follow recent motion customization methods (zhao2023motiondirector) with test-time fine-tuning for learning motion priors. Typically, reference videos are used to learn an appearance-debiased cross-video temporal LoRAs by injecting it into temporal attention layers with video-specific spatial LoRAs into spatial layers (wang2023modelscope; Zeroscope). Only temporal LoRAs are used during inference. Since our approach is based on CogVideoX (yang2024cogvideox), which employs 3D full attention instead of separate spatial-temporal attention, we manually designate even layers as “spatial” and odd layers as “temporal” to separate the learning of spatial and temporal LoRAs. We train LoRAs on top-ranked retrieved videos with all backbone parameters frozen, optimizing both diffusion loss for frame reconstruction and an appearance-debiased temporal loss to focus on motion-specific learning. Note that unlike previous methods using a single-prompt condition for all the retrieved videos, we utilize video caption provided from the database as per-video prompt. This helps the model implicitly separate motion-unrelated backgrounds, appearances, etc, allowing it to focus on motion-specific patterns. More design details are in Appendix A.2.
@@
VideoDirectorGPT (lin2023videodirectorgpt) 54.3 9.5 23.7 21.7 22.4 22.5 63.5 42.3 60.3 94.3
VLogger (zhuang2024vlogger) 62.5 41.3 23.5 23.1 22.5 22.2 73.6 43.4 61.2 96.2
DreamRunner (Ours)
70.7
(+13.1%)
55.1
(+33.4%)
24.7
(+5.11%)
23.7
(+2.60%)
24.2
(+7.56%)
24.1
(+8.56%)
93.6
(+27.2%)
55.4
(+27.6%)
62.1
(+1.47%)
98.1
(+1.98%)
Table 1: Evaluation of story-to-video generation on DreamStorySet. We compare ours with VideoDirectorGPT and VLogger on character consistency (CLIP and DINO scores), text instructions following and full prompt adherence (CLIP and ViCLIP scores), and event transitions smoothness (DINO score). Our relative improvement over VLogger is highlighted in blue.